MILLIMETER WAVE SPECTRUM OF THE NH_{3} - N_{2} COMPLEX

\underline{KALEY A. WALKER}^{a} and A. R. W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The only previous spectroscopic study of the ammonia - nitrogen van der Waals complex was made by Fraser et al.^{b} in 1986. They observed one main K = 0 series of rotational transitions, R(0) to R(3), in the 6 to 27 GHz region using the MBER technique. Another pair of lines near 13.5 GHz (reported in a footnote) was evidently the first member, R(1), of a series with K = 1. We have studied the NH_{3} - N_{2} spectrum in the 76 to 106 GHz region using a newly-developed pulsed jet millimeter wave spectrometer.^{c} Thirty-five lines were observed and assigned as R(11) to R(15) pure rotational transitions, belonging to six series which involve all combinations of the parent molecule nuclear spin species, ortho/para ammonia and ortho/para nitrogen. Two of these series are continuations to higher J-values of the transitions observed in the microwave region,^{b} and the remaining four series are new. We are thus able to predict a number of new microwave transitions for NH_{3} - N_{2} whose observation will help to confirm the present assignments of K-values and spin species.

aPresent address: Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada

bG.T. Fraser, D.D. Nelson, Jr., K.I. Peterson, and W. Klemperer, J. Chem. Phys. 84, 2472 (1986)

cK.A. Walker and A.R.W. McKellar, J. Mol. Spectrosc. 205, 331 (2001).